OpenStudy (anonymous):

3x3y + 12xy - 9x2y - 36y Part A: Rewrite the expression so that the GCF is factored completely. Show the steps of your work. (3 points) Part B: Rewrite the expression completely factored. Show the steps of your work. (4 points) Part C: If the two middle terms were switched so that the expression became 3x3y - 9x2y + 12xy - 36y, would the factored expression no longer be equivalent to your answer in part B? Explain your reasoning. (3 points)

OpenStudy (camerondoherty):

i dunno sry...

OpenStudy (anonymous):

3x^3y + 12xy - 9x^2y - 36y

OpenStudy (anonymous):

@jdoe0001 @dan815 @easyas123 can you guys help llovereading plz

OpenStudy (jdoe0001):

any ideas on the GCF?

OpenStudy (anonymous):

nope. im done for the day. my mind is literally tired. cant even think straight

OpenStudy (jdoe0001):

hhhehe... I meant @ilovehim121511 =)

OpenStudy (anonymous):

i would but i dont have time... sorry! DX

OpenStudy (jdoe0001):

ohh shoot.. wrong nick =)

OpenStudy (jdoe0001):

heheh ohh man I meant @Ilovereading

OpenStudy (jdoe0001):

so anyhow... \(\bf 3x^3y + 12xy - 9x^2y - 36y\implies {\color{brown}{ 3}}x^3{\color{brown}{ y}} + {\color{brown}{ 3}}\cdot 4x{\color{brown}{ y}} - {\color{brown}{ 3}}\cdot 3x^2{\color{brown}{ y}} - {\color{brown}{ 3}}\cdot 12{\color{brown}{ y}}\) what do you think @Ilovereading could be the GCF?

OpenStudy (jdoe0001):

ok,

OpenStudy (jdoe0001):

\(\bf 3x^3y + 12xy - 9x^2y - 36y\implies 3y(x^2+4x-3x^2-4y)\) right?

OpenStudy (anonymous):

so it would be 3y(x^3 + 4x -3x^2 -12)?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

hm actually shoot, got a typo should be \(\large \bf 3x^3y + 12xy - 9x^2y - 36y\implies 3y(x^2+4x-3x^2-12y)\)

OpenStudy (jdoe0001):

notice after plucking out the "3y" from the terms, all that's left is \(\large \bf {\color{brown}{ 3}}x^3{\color{brown}{ y}} + {\color{brown}{ 3}}\cdot 4x{\color{brown}{ y}} - {\color{brown}{ 3}}\cdot 3x^2{\color{brown}{ y}} - {\color{brown}{ 3}}\cdot 12{\color{brown}{ y}}\) is what's black

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so it would be (x^2+4)(x-3)?

OpenStudy (jdoe0001):

so to factor it... then well close

OpenStudy (jdoe0001):

hmmm shoot.. just notice yet another typoe... lemme correct that first \(\large \bf 3x^3y + 12xy - 9x^2y - 36y\implies 3y(x^2+4x-3x^2-12)\)

OpenStudy (anonymous):

then it would be 3y(x(x^2+4) - 3x^2 - 12 3y(x(x^2+4) - 3(x^2 + 4))?

OpenStudy (jdoe0001):

hmm one sec shoot lemem correct even that ohh may should be a 3 exponent there \(\large \bf 3x^3y + 12xy - 9x^2y - 36y\implies 3y(x^3+4x-3x^2-12)\)

OpenStudy (anonymous):

ok

OpenStudy (jdoe0001):

\(\bf 3x^3y + 12xy - 9x^2y - 36y\implies \begin{array}{llll} 3y[&x^3+4x&-3x^2-12]\\ &x(x^2+4)&-3(x^2+4)\\ &(x-3)&(x^2+4) \end{array}\)

OpenStudy (jdoe0001):

or \(\large \bf 3y[(x-3)(x^2+4)]\)

OpenStudy (anonymous):

oh okay so i was right. ok thanks! :) what about part C? i =m having trouble with that

OpenStudy (jdoe0001):

part C if we swap the middle terms ok so let's see that \(\bf 3x^3y {\color{brown}{ - 9x^2y+ 12xy}} - 36y\implies \begin{array}{llll} 3y[&x^3-3x^2&+4x-12]\\ &x^2(x-3)&+4(x-3)\\ &(x^2+4)& (x-3) \end{array}\\ 3y[(x^2+4)(x-3)]\)

OpenStudy (jdoe0001):

notice the two factored forms \(\large \bf { B\to 3y[(x-3){\color{brown}{ (x^2+4)}}]\\ C\to 3y[{\color{brown}{ (x^2+4)}}(x-3)] }\)

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

they're pretty much equivalent, since for factors, it doesn't quite matter the order 2* 4 = 8 or 4* 2 = 8 anyway

OpenStudy (anonymous):

 would i write that on my paper? the part with the question mark